OpenStudy (anonymous):

The sum of three consecutive integers is 267. What is the largest integer?

OpenStudy (anonymous):

90

OpenStudy (anonymous):

lets call the first integer (n), since they are consecutive, the other integers will be n+1 and n+2 so: n + (n+1) + (n+2) = 267 n + n +1 + n + 2 = 267 3n + 1 + 2 = 267 3n + 3 = 267 3n = 264 n = 88 so the numbers are just 88, 89, 90. The largest integer is 90.

OpenStudy (anonymous):

x+x+1+x+2=267 , solve the equation , and evaluate the x in x+2

OpenStudy (anonymous):

thankyou everyoneeee

OpenStudy (anonymous):

If x represents the total number of bags a person checks and each person checks more than two bags, the equation below can be used to determine C, the total fee for all the bags. 20 + 25 + 50(x - 2) = C Arianna paid a total of $145 to check all her bags on a flight. How many bags did she check in all?

OpenStudy (anonymous):

anyone know ?